OpenStudy (anonymous):

Which is a value of x, when 3x2 + 4 = 40? A) x = 2square root of 3 B) x = 2 square root of 6 C) x = 2 square root of 11 D) x = 3 square root of 2

OpenStudy (campbell_st):

subtract 4 from both sides \[3x^2 = 36\] divide by 3\[x^2 = 12\] take the square root of both sides \[x= \pm \sqrt{12}\] simplify the surd \[x = \pm 2\sqrt{3}\]

OpenStudy (anonymous):

thanks!!